given (x^2)/x^2-4

Find all critical values for 𝑔(π‘₯). Show all work and explain




b. Find 𝑔 β€²β€²(π‘₯). Then, use your answers from part (a) and the 2nd derivative test to determine if each critical value represents a relative maximum, minimum, or neither. Show all work and explain




c. On what interval(s), if any, is 𝑔(π‘₯) concave down. Show all work and explain


1
Expert's answer
2021-02-03T01:20:10-0500

"g(x) = \cfrac{x^2}{x^2-4}"


a) Find all critical values:

Critical values is values where first derivative is equal to zero ("g\prime(x)= 0" )

So find firs derivative:

"g\prime(x) = (\cfrac{x^2}{x^2-4})\prime = \cfrac{2x(x^2-4)-2x^3}{(x^2-4)^2}= -\cfrac{8x}{(x^2-4)^2}"

Find critical values:

"g\prime(x) = 0\\
-\cfrac{8x}{(x^2-4)^2} = 0 => x_0= 0"

So we have one critical value.


b) Relative maximum or minimum:

If "g\prime\prime(x_0) > 0," then "x_0" - is relative minimum

if "g\prime\prime(x_0) < 0," then "x_0" - is relative maximum.

So find "g\prime\prime(x)" :

"g\prime\prime(x) = \cfrac{-8(x^2-4)^2 + 32x^2(x^2-4)}{(x^2-4)^4} = \cfrac{-8(x^2-4)+32x}{(x^2-4)^3}"

"g\prime\prime(x_0) = -\cfrac{1}{2} < 0", so "x_0" is relative maximum


c) Where "g(x)" is concave down:

For all "x \in (-\infin, -2) \cup(2,+\infin)" "g\prime\prime(x) > 0", so on this intervals function is concave down.